# Constraints on Null Energy Condition Violation from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

### v1

@@ 1 Introduction
The detection of gravitational waves (GWs) by Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos has marked the beginning of a new era in observational astronomy LIGOScientific:2021djp . The ability to detect GWs has opened up a new window for studying the Universe, providing invaluable insights into the physics of compact objects, such as black holes and neutron stars. In addition to individual, loud GW events from compact binary coalescences, the superposition of a large number of weaker, unresolved GW signals can form a stochastic gravitational-wave background (SGWB). SGWBs can carry valuable information about the properties and distribution of their sources, both astrophysical and cosmological.
@@
Astrophysical sources of the SGWB include compact binary coalescences Zhu:2011bd ; Zhu:2012xw , and core-collapse supernovae Crocker:2017agi . Cosmological sources, on the other hand, are related to various physical processes in the early Universe, such as cosmic phase transitions Kibble:1980mv ; Witten:1984rs ; Mazumdar:2018dfl , primordial density perturbations during inflation Starobinsky:1979ty ; Turner:1996ck ; Bar-Kana:1994nri , cosmic strings Kibble:1976sj ; Sarangi:2002yt ; Damour:2004kw ; Siemens:2006yp ; LIGOScientific:2021nrg , and cosmic domain walls Vilenkin:1982ks ; Sikivie:1982qv . The detection of an SGWB of cosmological origin would provide unprecedented insights into the physics of the early Universe, potentially revealing new physics beyond the Standard Model.
@@
Although the LIGO-Virgo Collaboration has not yet detected a stochastic gravitational-wave background (SGWB) signal during their first three observing runs, they have established upper limits on its amplitude KAGRA:2021kbb . These constraints are instrumental in testing the viability of various cosmological models that predict an amplified primordial GW spectrum. Notably, a class of such models postulates a violation of the null energy condition (NEC) during the inflationary epoch of the early Universe Rubakov:2014jja ; Cai:2016thi ; Creminelli:2016zwa ; Cai:2017tku .
@@
The NEC is a fundamental principle in general relativity, stating that for any null vector, the contraction with the energy-momentum tensor yields a non-negative value. Nevertheless, scenarios that involve NEC violation have been conjectured to account for a range of early Universe phenomena, including the inflationary epoch and the genesis of primordial black holes Cai:2017dyi ; Kolevatov:2017voe ; Ilyas:2020zcb ; Ilyas:2020qja ; Zhu:2021ggm ; Cai:2023uhc ; see, for example, the review by Rubakov:2014jja . Advances in modified gravity theories, especially within the framework of “beyond Horndeski” models Libanov:2016kfc ; Kobayashi:2016xpl ; Ijjas:2016vtq ; Dobre:2017pnt ; Zhu:2021whu ; Cai:2022ori , have shown that stable NEC violations are theoretically feasible, circumventing the issue of pathological instabilities.
@@
The violation of the NEC has profound implications for the generation of primordial GWs in the early Universe. It has been shown that NEC violation can result in a notable amplification of the primordial tensor power spectrum at certain scales Cai:2020qpu ; Cai:2022nqv . This amplification manifests as a blue tilt in the tensor power spectrum, indicating an enhanced production of high-frequency GWs. As a consequence, the SGWB is expected to demonstrate a distinctive pattern that reflects the NEC-violating phase of the early Universe.
@@
In this paper, we investigate the constraints on NEC-violating inflationary models with GW data from Advanced LIGO and Advanced Virgo’s first three observing runs. We focus on a specific scenario that incorporates an intermediate NEC-violating phase during inflation proposed in Cai:2020qpu . This model predicts a characteristic blue-tilted tensor power spectrum at scales corresponding to the NEC-violating phase, which can potentially be probed by current and future GW observations. The rest of this paper is structured as follows. In Section 2, we begin by reviewing the key features of the inflationary model with an intermediate NEC-violating phase, focusing on its impact on the primordial tensor power spectrum. In Section 3, we describe the methodology used to derive constraints on the model parameters using the data from Advanced LIGO and Advanced Virgo. Lastly, we summarise our results and give a discussion of the implications in Section 4.
@@ 2 SGWB from NEC violation during inflation
In this section, we will provide a concise overview of the intermediate NEC violation during inflation. The scenario proposed by Cai:2020qpu describes the Universe starting with an initial phase of slow-roll inflation characterized by a Hubble parameter $H=H_{\text{inf1}}$. This is followed by a transition to a second phase of slow-roll inflation with a significantly larger Hubble parameter, $H=H_{\text{inf2}}$, after an intermediate stage of NEC violation. Throughout this evolution, the comoving Hubble horizon ($a^{-1}H^{-1}$) decreases, indicating the exit of perturbation modes from the horizon as the Universe undergoes accelerated expansion. These modes remain in the super-horizon regime until re-entering the horizon during the subsequent radiation-dominated or matter-dominated era.
@@
Perturbation modes exiting the horizon during the first ($k<k_{1}$) and second ($k>k_{2}$) stages of slow-roll inflation have nearly scale-invariant power spectra, consistent with CMB temperature anisotropy observations. However, the power spectrum amplitude of modes exiting during the second stage is significantly larger. The scale-invariance of the tensor power spectrum at large scales ensures a non-suppressed tensor-to-scalar ratio $r$ and slow-roll parameter in canonical single-field slow-roll inflation. At small scales, the scale-invariance of both scalar and tensor power spectra prevents them from growing to $\mathcal{O}(1)$, preserving the validity of perturbation theory at higher frequencies.
@@
The intermediate NEC violation during inflation can be realized with the effective field theory action Cai:2016thi :
$S=\int d^{4}x\sqrt{-g}\left[\frac{M_{P}^{2}}{2}R-\Lambda(t)-c(t)g^{00}+\frac{M_{2}^{4}(t)}{2}\left(\delta g^{00}\right)^{2}-\frac{m_{3}^{3}(t)}{2}\delta K\delta g^{00}+\frac{\tilde{m}_{4}^{2}(t)}{2}R^{(3)}\delta g^{00}\right].$ (1)
Here $\delta g^{00}=g^{00}+1$, $R^{(3)}$ denotes the Ricci scalar on the three-dimensional space-like hyper-surface, $\delta K=K-3H$, and $K$ is the extrinsic curvature. The time-dependent functions $c(t)$ and $\Lambda(t)$ determine the background evolution through the relations $c(t)=-M_{\rm P}^{2}\dot{H}$ and $\Lambda(t)=M_{\rm P}^{2}(\dot{H}+3H^{2})$. The functions $M_{2}^{4}(t)$, $m_{3}^{3}(t)$, and $\tilde{m}_{4}^{2}(t)$ are determined or constrained to ensure that the scalar perturbations agree with observations.
@@
Following Ye:2023tpz , we propose a parametrization of the primordial tensor power spectrum $P_{T}$ that captures the essential features:
$P_{T}=P_{T,1}+\frac{\pi}{4}(2-n_{T})\frac{k}{k_{2}}|g(k)|^{2}P_{T,2}~{},$ (2)
@@
$g(k)=H_{\frac{3-n_{T}}{2}}^{(1)}\left[\frac{2-n_{T}}{2}\frac{k}{k_{2}}\right]\sin\frac{k}{k_{2}}+H_{\frac{1-n_{T}}{2}}^{(1)}\left[\frac{2-n_{T}}{2}\frac{k}{k_{2}}\right]\left(\cos\frac{k}{k_{2}}-\frac{k_{2}}{k}\sin\frac{k}{k_{2}}\right)~{},$ (3)
@@
  • For modes with wavenumbers $k\ll k_{1}$, which exit the horizon during the first slow-roll inflationary stage, the second term in Eq. (2) becomes negligible compared to $P_{T,1}$, and thus $P_{T}\approx P_{T,1}$. This ensures that the power spectrum at large scales is nearly scale-invariant and consistent with CMB observations.
  • For modes with wavenumbers $k\gg k_{1}$, which exit the horizon during the second slow-roll inflationary stage or the NEC-violating stage, the second term in Eq. (2) dominates. This term captures the blue tilt and oscillatory features in the power spectrum around the transition scale $k\simeq k_{2}$, which are crucial for the generation of primordial black holes and the associated GW signals.
@@
$\Omega_{\mathrm{GW}}(f)=\frac{f}{\rho_{c}}\frac{\mathrm{d}\rho_{\mathrm{GW}}}{\mathrm{d}f},$ (4)
where $\rho_{c}=3H_{0}^{2}c^{2}/(8\pi G)$, with $c$ being the speed of light and $G$ representing Newton’s constant. The Hubble constant is taken from Planck 2018 observations as $H_{0}=67.4\mathrm{~{}km}\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ Planck:2018vyg . To connect the primordial tensor power spectrum $P_{T}$ with the observed GW energy spectrum $\Omega_{\rm{GW}}$, we employ the following relation Turner:1993vb :
$\Omega_{\rm{GW}}$ $=\frac{k^{2}}{12H_{0}^{2}}\left[\frac{3\Omega_{m}j_{l}(k\tau_{0})}{k\tau_{0}}\sqrt{1.0+1.36\frac{k}{k_{eq}}+2.50\left(\frac{k}{k_{eq}}\right)^{2}}\right]^{2}P_{T}$ (5)
$\simeq\frac{1}{24}\left(\frac{k}{H_{0}}\right)^{2}\left[\frac{3\Omega_{m}}{(k\tau_{0})^{2}}\sqrt{1.0+1.36\frac{k}{k_{eq}}+2.50\left(\frac{k}{k_{eq}}\right)^{2}}\right]^{2}P_{T}$
[…]
@@ 3 Data analysis
In this section, we describe the methodology employed to constrain the parameterized primordial tensor power spectrum, given by Eq. (2), using the data from Advanced LIGO and Virgo’s first three observing runs. As discussed in the previous section, we focus on the regime where $k_{1}\ll k$, in which the first term in Eq. (2) becomes negligible compared to the second term. This simplification allows us to safely neglect the contribution from $P_{T,1}$ and concentrate on the features arising from the NEC-violating stage and the second inflationary stage. With this approximation, the theoretical spectrum given by Eq. (2) can be fully characterized by three key parameters: $n_{T}$, $f_{c}$, and $P_{T,2}$.
To analyze the SGWB, we consider the LIGO-Hanford, LIGO-Livingston, and Virgo (HLV) network, with each detector labelled by the index $I={H,L,V}$. The time-series output of the detectors is denoted by $s_{I}(t)$, and its Fourier transform by $\tilde{s}_{I}(f)$. Following the formalism described in Romano:2016dpx ; Allen:1997ad , we define the cross-correlation statistic for the baseline $IJ$ as
$\hat{C}^{IJ}(f)=\frac{2}{T}\frac{\operatorname{Re}[\tilde{s}I^{\star}(f)\tilde{s}J(f)]}{\gamma{IJ}(f)S_{0}(f)},$ (6)
where $T$ represents the observation time, $\gamma{IJ}(f)$ is the normalized overlap reduction function Allen:1997ad for the baseline $IJ$, and the function $S_{0}(f)$ is defined as $S_{0}(f)=(3H_{0}^{2})/(10\pi^{2}f^{3})$. The estimator is normalized such that $\langle\hat{C}^{IJ}(f)\rangle=\Omega_{\mathrm{GW}}(f)$ in the absence of correlated noise. In the limit of a small signal-to-noise ratio, the variance can be approximated as
$\sigma_{IJ}^{2}(f)\approx\frac{1}{2T\Delta f}\frac{P_{I}(f)P_{J}(f)}{\gamma_{IJ}^{2}(f)S_{0}^{2}(f)},$ (7)
@@
We perform a Bayesian analysis to search for the SGWB from the NEC-violating phase during inflation using data from the Advanced LIGO and Advanced Virgo’s first three observing runs. The $\hat{C}^{IJ}(f)$ data, which is model-independent and publicly available KAGRA:2021kbb , is utilized in this analysis. To estimate the parameters of a specific SGWB model, we combine the spectra from each baseline $IJ$ to form the likelihood Mandic:2012pj
$p(\hat{C}^{IJ}(f_{k})|\boldsymbol{\theta})\propto\exp\left[-\frac{1}{2}\sum_{IJ}\sum_{k}\left(\frac{\hat{C}^{IJ}(f_{k})-\Omega_{\mathrm{M}}(f_{k}|\boldsymbol{\theta})}{\sigma_{IJ}^{2}(f_{k})}\right)^{2}\right],$ (8)
@@
$\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}=\frac{p(\hat{C}_{IJ}|\text{Model ofsignal})}{p(\hat{C}_{IJ}|\text{Pure noise})}=\frac{\int p(\hat{C}_{IJ}|\boldsymbol{\theta}_{\mathrm{GW}})\,p(\boldsymbol{\theta}_{\mathrm{GW}})\,\mathrm{d}\boldsymbol{\theta}_{\mathrm{GW}}}{\mathcal{N}}.$ (9)
Here, $\mathcal{N}$ is given by evaluating Eq. (8) for $\Omega_{\mathrm{M}}(f)=0$. When $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}>1$, there is support for the GW model compared to the NOISE model. The free parameters in the analysis are $\boldsymbol{\theta}_{\mathrm{GW}}\equiv(n_{T},f_{c},P_{T,2})$, and their priors are listed in Table 1. The Bayesian analysis is performed using the bilby package Ashton:2018jfp ; Romero-Shaw:2020owr and the dynamic nested sampling package dynesty Speagle:2019ivv .

### v2

# Constraints on Inflation with Null Energy Condition Violation from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

@@ 1 Introduction
The detection of gravitational waves (GWs) by Advanced LIGO [1] and Advanced Virgo [2] has marked the beginning of a new era in observational astronomy [3]. The ability to detect GWs has opened up a new window for studying the Universe, providing invaluable insights into the physics of compact objects, such as black holes and neutron stars. In addition to individual, loud GW events from compact binary coalescences, the superposition of a large number of weaker, unresolved GW signals can form a stochastic gravitational-wave background (SGWB). SGWBs can carry valuable information about the properties and distribution of their sources, both astrophysical and cosmological.
@@
Astrophysical sources of the SGWB include compact binary coalescences [4, 5], and core-collapse supernovae [6]. Cosmological sources, on the other hand, are related to various physical processes in the early Universe, such as cosmic phase transitions [7, 8, 9], primordial density perturbations during inflation [10, 11, 12], cosmic strings [13, 14, 15, 16, 17], and cosmic domain walls [18, 19]. The detection of an SGWB of cosmological origin would provide unprecedented insights into the physics of the early Universe, potentially revealing new physics beyond the Standard Model.
@@
Although the LIGO-Virgo Collaboration has not yet detected a stochastic gravitational-wave background (SGWB) signal during their first three observing runs, they have established upper limits on its amplitude [20]. These constraints are instrumental in testing the viability of various cosmological models that predict an amplified primordial GW spectrum. Notably, a class of such models postulates a violation of the null energy condition (NEC) during the inflationary epoch of the early Universe [21, 22, 23, 24].
@@
The NEC is a fundamental principle in general relativity, stating that for any null vector, the contraction with the energy-momentum tensor yields a non-negative value. Nevertheless, scenarios that involve NEC violation have been conjectured to account for a range of early Universe phenomena, including the inflationary epoch and the genesis of primordial black holes [25, 26, 27, 28, 29, 30, 31, 32]; see, for example, the review by [21]. Advances in modified gravity theories, especially within the framework of “beyond Horndeski” models [33, 34, 35, 36, 37, 38, 39], have shown that stable NEC violations are theoretically feasible, circumventing the issue of pathological instabilities.
@@
The violation of the NEC has profound implications for the generation of primordial GWs in the early Universe. It has been shown that NEC violation can result in a notable amplification of the primordial tensor power spectrum at certain scales [40, 41]. This amplification manifests as a blue tilt in the tensor power spectrum, indicating an enhanced production of high-frequency GWs. As a consequence, the SGWB is expected to demonstrate a distinctive pattern that reflects the NEC-violating phase of the early Universe.
@@
In this paper, we investigate the constraints on NEC-violating inflationary models with GW data from Advanced LIGO and Advanced Virgo’s first three observing runs. We focus on a specific scenario that incorporates an intermediate NEC-violating phase during inflation proposed in [40]. This model predicts a characteristic blue-tilted tensor power spectrum at scales corresponding to the NEC-violating phase, which can potentially be probed by current and future GW observations. The rest of this paper is structured as follows. In Section 2, we begin by reviewing the key features of the inflationary model with an intermediate NEC-violating phase, focusing on its impact on the primordial tensor power spectrum. In Section 3, we describe the methodology used to derive constraints on the model parameters using the data from Advanced LIGO and Advanced Virgo. Lastly, we summarise our results and give a discussion of the implications in Section 4.
@@ 2 SGWB from NEC violation during inflation
In this section, we will provide a concise overview of the intermediate NEC violation during inflation. The scenario proposed by [40] describes the Universe starting with an initial phase of slow-roll inflation characterized by a Hubble parameter $H=H_{\text{inf1}}$. This is followed by a transition to a second phase of slow-roll inflation with a significantly larger Hubble parameter, $H=H_{\text{inf2}}$, after an intermediate stage of NEC violation. Throughout this evolution, the comoving Hubble horizon ($a^{-1}H^{-1}$) decreases, indicating the exit of perturbation modes from the horizon as the Universe undergoes accelerated expansion. These modes remain in the super-horizon regime until re-entering the horizon during the subsequent radiation-dominated or matter-dominated era.
@@
Perturbation modes exiting the horizon during the first ($k<k_{1}$) and second ($k>k_{2}$) stages of slow-roll inflation have nearly scale-invariant power spectra, consistent with CMB temperature anisotropy observations. However, the power spectrum amplitude of modes exiting during the second stage is significantly larger. The scale-invariance of the tensor power spectrum at large scales ensures a non-suppressed tensor-to-scalar ratio $r$ and slow-roll parameter in canonical single-field slow-roll inflation. At small scales, the scale-invariance of both scalar and tensor power spectra prevents them from growing to $\mathcal{O}(1)$, preserving the validity of perturbation theory at higher frequencies. In the specific scenario we have considered, the NEC violation occurs during an intermediate stage between two slow-roll inflationary phases. The end of inflation is attributed to the dynamics of the second slow-roll phase, which is assumed to follow the standard mechanism for ending inflation, such as the inflaton field reaching the minimum of its potential.
@@
The intermediate NEC violation during inflation can be realized with the effective field theory action [22]:
$S=\int d^{4}x\sqrt{-g}\left[\frac{M_{P}^{2}}{2}R-\Lambda(t)-c(t)g^{00}+\frac{M_{2}^{4}(t)}{2}\left(\delta g^{00}\right)^{2}-\frac{m_{3}^{3}(t)}{2}\delta K\delta g^{00}+\frac{\tilde{m}_{4}^{2}(t)}{2}R^{(3)}\delta g^{00}\right].$ (2.1)
Here $\delta g^{00}=g^{00}+1$, $R^{(3)}$ denotes the Ricci scalar on the three-dimensional space-like hyper-surface, $\delta K=K-3H$, and $K$ is the extrinsic curvature. The time-dependent functions $c(t)$ and $\Lambda(t)$ determine the background evolution through the relations $c(t)=-M_{\rm P}^{2}\dot{H}$ and $\Lambda(t)=M_{\rm P}^{2}(\dot{H}+3H^{2})$. The functions $M_{2}^{4}(t)$, $m_{3}^{3}(t)$, and $\tilde{m}_{4}^{2}(t)$ are determined or constrained to ensure that the scalar perturbations agree with observations. The EFT approach allows us to systematically incorporate modifications to general relativity and study their phenomenological consequences without committing to a specific underlying theory of quantum gravity. In the context of NEC violation, the EFT framework can accommodate such features by including higher-order operators that modify the standard kinetic and potential terms of the scalar field. These modifications can lead to scenarios where the usual energy conditions are not satisfied. Specifically, terms in the action such as $\frac{\tilde{m}_{4}^{2}(t)}{2}R^{(3)}\delta g^{00}$ and higher derivative interactions can effectively lead to negative contributions to the energy density, thus violating the NEC [22].
@@
Following [42], we propose a parametrization of the primordial tensor power spectrum $P_{T}$ that captures the essential features:
$P_{T}=P_{T,1}+\frac{\pi}{4}(2-n_{T})\frac{k}{k_{2}}|g(k)|^{2}P_{T,2}~{},$ (2.2)
@@
$g(k)=H_{\frac{3-n_{T}}{2}}^{(1)}\left[\frac{2-n_{T}}{2}\frac{k}{k_{2}}\right]\sin\frac{k}{k_{2}}+H_{\frac{1-n_{T}}{2}}^{(1)}\left[\frac{2-n_{T}}{2}\frac{k}{k_{2}}\right]\left(\cos\frac{k}{k_{2}}-\frac{k_{2}}{k}\sin\frac{k}{k_{2}}\right)~{},$ (2.3)
@@
  • For modes with wavenumbers $k\ll k_{1}$, which exit the horizon during the first slow-roll inflationary stage, the second term in Eq. (2.2) becomes negligible compared to $P_{T,1}$, and thus $P_{T}\approx P_{T,1}$. This ensures that the power spectrum at large scales is nearly scale-invariant and consistent with CMB observations.
  • For modes with wavenumbers $k\gg k_{1}$, which exit the horizon during the second slow-roll inflationary stage or the NEC-violating stage, the second term in Eq. (2.2) dominates. This term captures the blue tilt and oscillatory features in the power spectrum around the transition scale $k\simeq k_{2}$, which are crucial for the generation of primordial black holes and the associated GW signals.
@@
$\Omega_{\mathrm{GW}}(f)=\frac{f}{\rho_{c}}\frac{\mathrm{d}\rho_{\mathrm{GW}}}{\mathrm{d}f},$ (2.4)
where $\rho_{c}=3H_{0}^{2}c^{2}/(8\pi G)$, with $c$ being the speed of light and $G$ representing Newton’s constant. The Hubble constant is taken from Planck 2018 observations as $H_{0}=67.4\mathrm{~{}km}\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ [43]. To connect the primordial tensor power spectrum $P_{T}$ with the observed GW energy spectrum $\Omega_{\rm{GW}}$, we employ the following relation [44]:
$\Omega_{\rm{GW}}$ $=\frac{k^{2}}{12H_{0}^{2}}\left[\frac{3\Omega_{m}j_{l}(k\tau_{0})}{k\tau_{0}}\sqrt{1.0+1.36\frac{k}{k_{\text{eq}}}+2.50\left(\frac{k}{k_{\text{eq}}}\right)^{2}}\right]^{2}P_{T}$ (2.5)
$\simeq\frac{1}{24}\left(\frac{k}{H_{0}}\right)^{2}\left[\frac{3\Omega_{m}}{(k\tau_{0})^{2}}\sqrt{1.0+1.36\frac{k}{k_{\text{eq}}}+2.50\left(\frac{k}{k_{\text{eq}}}\right)^{2}}\right]^{2}P_{T}$
where $\tau_{0}=1.41\times 10^{4}$ Mpc is the conformal time today, $k_{\text{eq}}=0.073\,\Omega_{\text{m}}h^{2}$ Mpc-1 is the wavenumber at matter-radiation equality, $\Omega_{\rm m}$ is the present matter density fraction, and the frequency $f$ is related to the wavenumber $k$ by $k=2\pi f$. This relation enables us to predict the observable GW energy spectrum based on our parametrization of the primordial tensor power spectrum, facilitating a direct comparison with current and future GW experiments.
@@ 3 Data analysis
In this section, we describe the methodology employed to constrain the parameterized primordial tensor power spectrum, given by Eq. (2.2), using the data from Advanced LIGO and Virgo’s first three observing runs. As discussed in the previous section, we focus on the regime where $k_{1}\ll k$, in which the first term in Eq. (2.2) becomes negligible compared to the second term. This simplification allows us to safely neglect the contribution from $P_{T,1}$ and concentrate on the features arising from the NEC-violating stage and the second inflationary stage. With this approximation, the theoretical spectrum given by Eq. (2.2) can be fully characterized by three key parameters: $n_{T}$, $f_{c}$, and $P_{T,2}$, where $f_{c}\equiv 2\pi k_{2}$.
To analyze the SGWB, we consider the LIGO-Hanford, LIGO-Livingston, and Virgo (HLV) network, with each detector labelled by the index $I={H,L,V}$. The time-series output of the detectors is denoted by $s_{I}(t)$, and its Fourier transform by $\tilde{s}_{I}(f)$. Following the formalism described in [45, 46], we define the cross-correlation statistic for the baseline $IJ$ as
$\hat{C}^{IJ}(f)=\frac{2}{T}\frac{\operatorname{Re}[\tilde{s}I^{\star}(f)\tilde{s}J(f)]}{\gamma{IJ}(f)S_{0}(f)},$ (3.1)
where $T$ represents the observation time, $\gamma{IJ}(f)$ is the normalized overlap reduction function [46] for the baseline $IJ$, and the function $S_{0}(f)$ is defined as $S_{0}(f)=(3H_{0}^{2})/(10\pi^{2}f^{3})$. The estimator is normalized such that $\langle\hat{C}^{IJ}(f)\rangle=\Omega_{\mathrm{GW}}(f)$ in the absence of correlated noise. In the limit of a small signal-to-noise ratio, the variance can be approximated as
$\sigma_{IJ}^{2}(f)\approx\frac{1}{2T\Delta f}\frac{P_{I}(f)P_{J}(f)}{\gamma_{IJ}^{2}(f)S_{0}^{2}(f)},$ (3.2)
@@
We perform a Bayesian analysis to search for the SGWB from the NEC-violating phase during inflation using data from the Advanced LIGO and Advanced Virgo’s first three observing runs. The $\hat{C}^{IJ}(f)$ data, which is model-independent and publicly available [20], is utilized in this analysis. To estimate the parameters of a specific SGWB model, we combine the spectra from each baseline $IJ$ to form the likelihood [47]
$p(\hat{C}^{IJ}(f_{k})|\boldsymbol{\theta})\propto\exp\left[-\frac{1}{2}\sum_{IJ}\sum_{k}\left(\frac{\hat{C}^{IJ}(f_{k})-\Omega_{\mathrm{M}}(f_{k}|\boldsymbol{\theta})}{\sigma_{IJ}^{2}(f_{k})}\right)^{2}\right],$ (3.3)
@@
$\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}=\frac{p(\hat{C}_{IJ}|\text{Model ofsignal})}{p(\hat{C}_{IJ}|\text{Pure noise})}=\frac{\int p(\hat{C}_{IJ}|\boldsymbol{\theta}_{\mathrm{GW}})\,p(\boldsymbol{\theta}_{\mathrm{GW}})\,\mathrm{d}\boldsymbol{\theta}_{\mathrm{GW}}}{\mathcal{N}}.$ (3.4)
Here, $\mathcal{N}$ is given by evaluating Eq. (3.3) for $\Omega_{\mathrm{M}}(f)=0$. When $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}>1$, there is support for the GW model compared to the NOISE model. The free parameters in the analysis are $\boldsymbol{\theta}_{\mathrm{GW}}\equiv(n_{T},f_{c},P_{T,2})$, and their priors are listed in Table 1. The Bayesian analysis is performed using the bilby package [48, 49] and the dynamic nested sampling package dynesty [50].